OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. 2x + y + z = –7 x – 3y + 4z = –14 x – 2y – 3z = –11

OpenStudy (across):

Matrices! That answer would blow your professor away.

OpenStudy (anonymous):

what hahahahaha

OpenStudy (zarkon):

I would use an augmented matrix and row reduce to get \[[x,y,z]=[-5,3,0]\]

OpenStudy (across):

I would code a genetic algorithm with fitness function \[f(x)=\frac{1}{[a(x)+7]^2+[b(x)+14]^2+[c(x)+11]^2}\]where \[a(x)=2x+y+z\]\[b(x)=x-3y+4z\]\[c(x)=x-2y-3z\]:P

OpenStudy (zarkon):

how about let \[f(x,y,z)=(2x + y + z +7)^2+ (x – 3y + 4+14)^2+(x – 2y – 3z+11)^2\] then use the steepest decent algorithm to find the minimum of f